OpenStudy (tlar):

Let f(x) = (1-4x^2)/(x) For what values of x is f decreasing? For what values of x is the graph of f concave down?

OpenStudy (tlar):

@Badsheep

OpenStudy (solomonzelman):

1. \(\color{black}{\displaystyle f(x) }\) increases on \(\color{black}{\displaystyle I }\), iff \(\color{black}{\displaystyle f'(x)>0 }\) on \(\color{black}{\displaystyle I }\). 2. \(\color{black}{\displaystyle f(x) }\) decreases on \(\color{black}{\displaystyle I }\), iff \(\color{black}{\displaystyle f'(x)<0 }\) on \(\color{black}{\displaystyle I }\).

OpenStudy (solomonzelman):

So, to find \(\color{black}{{\rm for~what~values~}f~{\rm is~decreasing}}\), all you need to do find \(\color{black}{\displaystyle f'(x) }\) and to then set \(\color{black}{\displaystyle f'(x)<0 }\) and solve for \(\color{black}{\displaystyle x }\).

OpenStudy (solomonzelman):

The relation b/w \(\color{black}{\displaystyle f }\) and \(\color{black}{\displaystyle f' }\), is the same relation as that b/w \(\color{black}{\displaystyle f' }\) and \(\color{black}{\displaystyle f''}\). In other words, 1b). \(\color{black}{\displaystyle f'(x) }\) (or the slope of \(\color{black}{\displaystyle f }\)) increases on \(\color{black}{\displaystyle I }\), iff \(\color{black}{\displaystyle f''(x)>0 }\) on \(\color{black}{\displaystyle I }\). 2b). \(\color{black}{\displaystyle f'(x) }\) (or the slope of \(\color{black}{\displaystyle f }\)) decreases on \(\color{black}{\displaystyle I }\), iff \(\color{black}{\displaystyle f''(x)<0 }\) on \(\color{black}{\displaystyle I }\).

OpenStudy (solomonzelman):

What "concave down" means, is that the (not the \(f\), but the) slope of \(f\) is decreasing. So, in order to find where \(f\) is concave down, you need to find \(f''(x)\) and then solve \(f''(x)<0\).